A Nissan Motor Corporation advertisement read, “The average man’s I.Q. is 107. The average brown trout’s I.Q. is 4. So why can’t man catch brown trout?” Suppose you believe that the brown trout’s mean I.Q. is greater than four. You catch 12 brown trout. A fish psychologist determines the I.Q.s as follows: 5; 4; 7; 3; 6; 4; 5; 3; 6; 3; 8; 5. Conduct a hypothesis test of your belief.
Answer by Boreal(15235)

Mean is 59/12=4.916667
s=1.62
Ho: IQ<=4
Ha:IQ>4
alpha=0.05 p{reject Ho|Ho true}
test statistic is t df=11 =(x bar-mean)/s/sqrt(n)
critical value t>1.796
calculation is t=(0.916667)/1.62/sqrt(12))=1.96
reject Ho and IQ >4
p-value 0.038